An annuity that pays out payments that increase by 0.75% each payment period will be paid out over a 10 year period will be purchased with interest of 6% compounded monthly. If the payments will start at $1000 the first month, how much will need to be in the account to sustain the 10 years of payments?
Answer #1

Calculating the amount needed to sustain the paymenst for 10 years:-

Present Value of Growing Annuity = C*\frac{1-[(1+g)^{n}*(1+r)^{-n}]}{r-g}

Where, C= First Payment= $1000

r = Periodic Interest rate = 6%/12 =0.5%

g = Growth rate of further payments = 0.75%

n= no of periods = 10years*12 = 120

Present Value = 1000*\frac{1-[(1+0.0075)^{120}*(1+0.005)^{-120}]}{(0.005-0.0075)}

Present Value = 1000*\frac{1-[2.45135707812*0.54963273336]}{(-0.0025)}

Present Value = $138,938.44

So, will need to be in the account to sustain the 10 years of payments is $138,938.44

If you need any clarification, you can ask in comments.     

If you like my answer, then please up-vote as it will be motivating